OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). A:y = 2x + 6 B:y =−1/2x + 3 C:y =−1/2x +6 D: y = 2x + 3

OpenStudy (kropot72):

The product of the slope of the given equation and the slope of the equation that is perpendicular must equal -1. Let the slope of the equation that is perpendicular = m 2 * m = -1 .............(1) As a first step in the solution, can you solve equation (1) ?

OpenStudy (anonymous):

O.o i'm cunfused :P

OpenStudy (anonymous):

Is the answer D? I'm not good with slopes

OpenStudy (kropot72):

To solve 2m = -1 divide each side of the equation by 2. \[\frac{2\times m}{2}=-\frac{1}{2}\] Can you see the value of m now?

OpenStudy (anonymous):

Still not getting it... Thanks for your time.